OpenStudy (anonymous):

The number of bowls of soup served was sampled during a lunch period 11:00 a.m. – 2:30 p.m. at a soup stand on a weekday. Below is the sample: 71 63 92 58 93 74 92 99 82 63 92 91 63 54 76 95 76 Construct a stem and leaf plot of the data.

OpenStudy (anonymous):

i really need help with this one, and my online tutors are not in yet. thank you so much

OpenStudy (anonymous):

Rearrange numbers in order 54 58 63 63 63 71 74 76 76 91 92 92 92 95 101 112 116 Mean: 54+58+63+63+63+71+74+76+76+91+92+92+92+9… = 1389 / 17 = 81.71 Median: 76 Mode: 63 & 92 Midrange: 54 + 116 = 170 / 2 = 85 The 1st quartile is the median of the numbers to the left of the true median. 1st quartile: 63 + 63 = 126 / 2 = 63 The 3rd quartile is the median of the numbers to the right of the true median. 3rd quartile: 92 + 95 = 187 / 2 = 93.5

OpenStudy (anonymous):

Thank you so much lego, I really appreciate it.

OpenStudy (anonymous):

Your a real life saver....